40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14251     Entry time: Sat Oct 13 20:11:10 2018     In reply to: 14248     Reply to this: 14254
Author: yuki 
Type: Update 
Category: Computer Scripts / Programs 
Subject: loss measurements 
Quote:

the script "armloss_AS_calc.py",

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.

Some changes were made in the script for getting the signals of beam power:

  • The script sees "C1:ASS-X(Y)ARM_ETM_PIT/YAW_L_DEMOD_I_OUTPUT" and stops running until the signals become small, however some offset could be on the signal. So I changed it into waiting until (DEMOD - OFFSET) becomes small. (Yesterday I wrote ETM_YAW_L_DEMOD_I_OUTPUT was about 15 and was little noisy. I was wrong. That was just a offset value.)
  • I added a code which stops running the script when the power of transmitted IR beam is low. You can set this threshold. The nominal value of "C1:LSC-TRX(Y)_OUT16" is 1.2 (1.0), so the threshold is set 0.8 now.  

In the yesterday measurement the beam power of ASDC is higher when locked than when misaligned and I wrote it maybe caused by over-coupled cavity. Then I did a calculation as following to explain this:

  • assume power transmissivity of ITM and ETM are 1.4e-2 and 1.4e-5.
  • assume loss-less mirror, you can calculate amplitude reflectivity of ITM and ETM.
  • consider a cavity which consists two mirrors and is loss-less, then \frac{E_{r}}{E_{in}} = \frac{-r_1+r_2e^{i\phi}}{1-r_1r_2e^{i\phi}} holds. r1 and r2 are amplitude reflectivity of ITM and ETM, and E is electric filed.
  • Then you can calculate the power of reflected beam when resonated and when anti-resonated. The fraction of these value is \frac{P_{RESONANT}}{P_{ANTI-RESO}} = 0.996, which is smaller than 1.
  • I found this calculation was wrong! Above calculatation only holds when cavity is aligned, not when misaligned. 99.04% of incident beam power reflects when locked, and (100-1.4)% reflects when misaligned. The proportion is P(locked)/P(misaligned)=1.004, higher than 1.

 

ELOG V3.1.3-